OpenStudy (anonymous):

given that tanθ=2/3 and θ is in the 3rd quadrant.without evaluating the angle θ,find the exact values of a)cos(-θ) b)sin(-θ) c)sin(θ-8π)

OpenStudy (anonymous):

My Mind has got rusted.. :)

OpenStudy (anonymous):

really thanks for the help

OpenStudy (anonymous):

Use : \[1 + \tan^2(\theta) = \sec^2(\theta)\] Here find sec(theta) first...

OpenStudy (anonymous):

But theta here is in third quadrant, so sec(theta) will be negative..

OpenStudy (anonymous):

\[\sec(\theta) = -\sqrt{\frac{5}{3}}\]

OpenStudy (anonymous):

Sorry that is not 4, that is 5 there..

OpenStudy (anonymous):

Just take the reciprocal of sec(theta) to find cos(theta) ie. \[\cos(\theta) = \frac{1}{\sec(\theta)}\] Now using : \[\sin(\theta) = \cos(\theta) \cdot \tan(\theta)\] Find sin(theta)..

OpenStudy (anonymous):

the ans for a is cos(-θ)=-3/Square root 13 the ans for b is sin(-θ)=2/Square root 13 the ans for c is sin(θ-8π)=-2/Square root 13 i got the answer but i not sure how to do

OpenStudy (kropot72):

@waterineyes \[\tan ^{2} (\theta)=(\frac{2}{3})^{2}=\frac{4}{9}\]

OpenStudy (anonymous):

I said my mind has got rusty.. :) @kropot72

OpenStudy (anonymous):

\[1 + (\frac{2}{3})^2 = \sec^2(\theta) \implies \sec(\theta) = 1 + \frac{4}{9} \implies \sec^2(\theta) = \frac{13}{9}\]

OpenStudy (anonymous):

Now find sed(theta) here.. \[\sec(\theta) = \sqrt{\frac{13}{9}} \implies \frac{\sqrt{13}}{3}\]

OpenStudy (kropot72):